# Generalized Moving Least-Squares for Solving Vector-valued PDEs on Unknown Manifolds

### v1

@@ 3.2.3 Matrix assembly of the Bochner Laplacian
In this section, we discuss how to calculate the weights $\{\bm{w}^{k}\}$ in (12) from the weights $\{\bm{\tilde{w}}^{k}\}$ and assemble all $\{\bm{w}^{k}\}$ into the appropriate rows and columns of a sparse $2N\times 2N$ Laplacian matrix $\mathbf{L}_{B}$. So far, we have obtained the components$\ \{\tilde{u}({\mathbf{x}_{0,k}})\in\mathbb{R}^{2\times 1}\}_{k=1}^{K}$ and the weights $\{\bm{\tilde{w}}^{k}\}_{k=1}^{K}$ to approximate $\Delta_{B}u({\mathbf{x}_{0}})$ (see equation (16)) in the local bases $\left\{(\partial_{1}|_{\mathbf{x}_{0,k}},\partial_{2}|_{\mathbf{x}_{0,k}})\right\}_{k=1}^{K}$ of each stencil $S_{{\mathbf{x}_{0}}}$. In order to construct a sparse $2N\times 2N$ Bochner Laplacian matrix, we need to transform the weights from the local bases $\left\{\partial_{1}|_{\mathbf{x}_{0,k}},\partial_{2}|_{\mathbf{x}_{0,k}}\right\}_{k=1}^{K}$ in each $S_{{\mathbf{x}_{0}}}$ to the unified global basis $\{\bm{t}_{1}(\mathbf{x}_{0,k}),\bm{t}_{2}(\mathbf{x}_{0,k})\}_{k=1}^{K}$.
@@
$\sum_{k=1}^{K}\bm{w}^{k}u({\mathbf{x}_{0,k}})=(\bm{w}^{1},\ldots,\bm{w}^{K})\mathbf{u}_{{\mathbf{x}_{0}}}\approx\Delta_{B}u({\mathbf{x}_{0}})\approx\sum_{k=1}^{K}\bm{\tilde{w}}^{k}\tilde{u}({\mathbf{x}_{0,k}}).$ (20)
@@
$\bm{w}^{k}u({\mathbf{x}_{0,k}})\approx\bm{\tilde{w}}^{k}\tilde{u}({\mathbf{x}_{0,k}}),\text{ \ \ }\forall k=1,\ldots,K.$
Besides, notice that $u=u^{j}\bm{t}_{j}\approx\tilde{u}^{i}\partial_{i}$ in each $S_{{\mathbf{x}_{0}}}$, that is,
$\left[\begin{matrix}\bm{t}_{1}({\mathbf{x}_{0,k}})&\bm{t}_{2}({\mathbf{x}_{0,k}})\end{matrix}\right]u({\mathbf{x}_{0,k}})\approx\left[\begin{matrix}\partial_{1}|_{{\mathbf{x}_{0,k}}}&\partial_{2}|_{{\mathbf{x}_{0,k}}}\end{matrix}\right]\tilde{u}({\mathbf{x}_{0,k}})\in\mathbb{R}^{3\times 1},$
@@
The complete procedure of the intrinsic GMLS approximation of the Bochner Laplacian is listed in Algorithm 1. Incidentally, one can follow the similar procedure to approximate other vector Laplacians acting on vector fields using the intrinsic method (See Appendix A).
Algorithm 1 Intrinsic GMLS of the Bochner Laplacian
@@
1:  Input: A point cloud $\{\mathbf{x}_{i}\}_{i=1}^{N}\subset M$, bases of (estimated) tangent vector space at each node $\left\{\bm{t}_{1}(\mathbf{x}_{i}),\bm{t}_{2}(\mathbf{x}_{i})\right\}_{i=1}^{N}$, the degree $l$ of polynomials, and a parameter $K>(l+2)(l+1)/2$ nearest neighbors.
2:  Set $\mathbf{L}_{B}$ to be a sparse $2N\times 2N$ matrix with $4NK$ nonzeros.
@@
5:     Find the GMLS approximation $\hat{q}$ of the local embedding map as in (9).
6:     Construct the local basis of $\{\partial_{1}|_{{\mathbf{x}_{i,k}}},\partial_{2}|_{{\mathbf{x}_{i,k}}}\}$ over the stencil $S_{\mathbf{x}_{i}}$ as in (10).
7:     Compute the geometric quantities in (15) using the coefficients of $\hat{q}$.
8:     Calculate the weights $\{\bm{\tilde{w}}^{k}\}_{k=1}^{K}$ using (19).
9:     Calculate the weights $\{\bm{w}^{k}\}_{k=1}^{K}$ as
$\bm{w}^{k}=\bm{\tilde{w}}^{k}\left(\left[\begin{matrix}\partial_{1}^{\top}|_{{\mathbf{x}_{0,k}}}\\ \partial_{2}^{\top}|_{{\mathbf{x}_{0,k}}}\end{matrix}\right]\left[\begin{matrix}\partial_{1}|_{{\mathbf{x}_{0,k}}}&\partial_{2}|_{{\mathbf{x}_{0,k}}}\end{matrix}\right]\right)^{-1}\left[\begin{matrix}\partial_{1}^{\top}|_{{\mathbf{x}_{0,k}}}\\ \partial_{2}^{\top}|_{{\mathbf{x}_{0,k}}}\end{matrix}\right]\left[\begin{matrix}\bm{t}_{1}({\mathbf{x}_{0,k}})&\bm{t}_{2}({\mathbf{\ x}_{0,k}})\end{matrix}\right].$
10:     Arrange the weights $\{\bm{w}^{k}\}_{k=1}^{K}$ into corresponding rows and columns of $\mathbf{L}_{B}$.
11:  end for
12:  Output: The approximate operator matrix $\mathbf{L}_{B}$.
@@ 3.3 Approximation of covariant derivative
Note that the covariant derivative of the vector field $u=\tilde{u}^{i}\partial_{i}$ along itself could be written as
$\nabla_{u}u=\tilde{u}_{;k}^{i}\tilde{u}^{k}\partial_{i}=\tilde{u}^{k}\left(\partial_{k}\tilde{u}^{i}+\tilde{u}^{j}\Gamma_{kj}^{i}\right)\partial_{i}.$
By using the local Monge coordinate system introduced in Section 3.1, the approximation of $\nabla_{u}u$ could be derived directly as
$\nabla_{u}u(\mathbf{x}_{0})$ $=$ $\tilde{u}^{k}\left(\partial_{k}\tilde{u}^{i}+\tilde{u}^{j}\Gamma_{kj}^{i}\right)\partial_{i}\big{|}_{\mathbf{x}_{0}}$
@@
where the coefficients $\beta_{\alpha_{1},\alpha_{2}}$ and $\gamma_{\alpha_{1},\alpha_{2}}$ for $0\leq\alpha_{1}+\alpha_{2}\leq l$ are defined in (17). Here, we again use Proposition 3.1 to simplify the calculations.

### v2

@@ 3.2.3 Matrix assembly of the Bochner Laplacian
In this section, we discuss how to calculate the weights $\{\bm{w}^{k}\}$ in (13) from the weights $\{\bm{\tilde{w}}^{k}\}$ in (17) and assemble all $\{\bm{w}^{k}\}$ into the appropriate rows and columns of a sparse $2N\times 2N$ Laplacian matrix $\mathbf{L}_{B}$. So far, we have obtained the components$\ \{\bm{\tilde{u}}|_{{\mathbf{x}_{0,k}}}\in\mathbb{R}^{2\times 1}\}_{k=1}^{K}$ and the weights $\{\bm{\tilde{w}}^{k}\}_{k=1}^{K}$ to approximate $\Delta_{B}\bm{u}({\mathbf{x}_{0}})$ (see equation (17)) in the local bases $\left\{(\partial_{1}|_{\mathbf{x}_{0,k}},\partial_{2}|_{\mathbf{x}_{0,k}})\right\}_{k=1}^{K}$ of each stencil $S_{{\mathbf{x}_{0}}}$. In order to construct a sparse $2N\times 2N$ Bochner Laplacian matrix, we need to transform the weights from the local bases $\left\{\partial_{1}|_{\mathbf{x}_{0,k}},\partial_{2}|_{\mathbf{x}_{0,k}}\right\}_{k=1}^{K}$ in each $S_{{\mathbf{x}_{0}}}$ to the unified global basis $\{\bm{t}_{1}(\mathbf{x}_{0,k}),\bm{t}_{2}(\mathbf{x}_{0,k})\}_{k=1}^{K}$.
@@
$\sum_{k=1}^{K}\bm{w}^{k}\bm{u}|_{{\mathbf{x}_{0,k}}}=(\bm{w}^{1},\ldots,\bm{w}^{K})\mathbf{u}_{{\mathbf{x}_{0}}}\approx\Delta_{B}\bm{u}({\mathbf{x}_{0}})\approx\sum_{k=1}^{K}\bm{\tilde{w}}^{k}\bm{\tilde{u}}|_{{\mathbf{x}_{0,k}}}.$ (21)
@@
$\bm{w}^{k}\bm{u}|_{{\mathbf{x}_{0,k}}}\approx\bm{\tilde{w}}^{k}\bm{\tilde{u}}|_{{\mathbf{x}_{0,k}}},\text{ \ \ }\forall k=1,\ldots,K.$
We note that $\bm{u}=u^{j}\bm{t}_{j}\approx\tilde{u}^{i}\partial_{i}$ in each $S_{{\mathbf{x}_{0}}}$, that is,
$\left[\begin{matrix}\bm{t}_{1}({\mathbf{x}_{0,k}})&\bm{t}_{2}({\mathbf{x}_{0,k}})\end{matrix}\right]\bm{u}|_{{\mathbf{x}_{0,k}}}\approx\left[\begin{matrix}\partial_{1}|_{{\mathbf{x}_{0,k}}}&\partial_{2}|_{{\mathbf{x}_{0,k}}}\end{matrix}\right]\bm{\tilde{u}}_{{\mathbf{x}_{0,k}}}\in\mathbb{R}^{3\times 1},$
@@
3.3 Intrinsic formulation in higher dimensional spaces
We now extend our approach to the general $d$-dimensional manifolds embedded in $n$-dimensional ambient spaces ($d\ll n$). In numerical implementation, we focus on small $d$ ranging around $1\sim 3$ and relatively large $n$ on the order of $O(10)$. In Section 3.3.1, we first derive the intrinsic formulation for the Bochner Laplacian in the local Monge parametrization. However, a direct implementation of the formulation in Section 3.3.1 results in a computationally expensive algorithm with a complexity of $O(Nn^{3})$ (due to the full QR factorization to generate normal directions). To reduce the complexity for manifold approximation, we introduce a practical algorithm in Section 3.3.2 that reduces the complexity to $O(Nn)$. In Section 3.3.3, we present the complete algorithm along with its complexity analysis.
3.3.1 Intrinsic formulation in the Monge parametrization
For the ease of discussion, we again use $\left\{\bm{t}_{i}({\mathbf{x}_{j}})\in\mathbb{R}^{n}\right\}_{i=1,\ldots,d}^{j=1,\ldots,N}$ to denote the global tangent vector basis on the point cloud $\left\{{\mathbf{x}_{j}}\right\}_{j=1}^{N}$ for both known and unknown manifolds. Then one can define a local coordinate chart for the manifold near the base point,
${\mathbf{x}}(\theta_{1},\ldots,\theta_{d})={\mathbf{x}_{0}}+\sum_{i=1}^{d}\theta_{i}\bm{t}_{i}({\mathbf{x}_{0}})+\sum_{s=1}^{n-d}q^{s}\left(\theta_{1},\ldots,\theta_{d}\right)\bm{n}_{s}({\mathbf{x}_{0}}),$ (22)
where $\theta_{i}\left({\mathbf{x}}\right)=\bm{t}_{i}({\mathbf{x}_{0}})\cdot(\mathbf{x}-{\mathbf{x}_{0}})$ and $q^{s}\left({\mathbf{x}}\right)=\bm{n}_{s}({\mathbf{x}_{0}})\cdot(\mathbf{x}-{\mathbf{x}_{0}}).$ The tangent vector bases in the Monge coordinate system are $\partial_{k}=\frac{\partial}{\partial\theta_{k}}=\bm{t}_{k}({\mathbf{x}_{0}})+\sum_{j=1}^{n-d}\frac{\partial q^{j}}{\partial\theta_{k}}\bm{n}_{j}({\mathbf{x}_{0}}),\text{ \ for }k=1,\ldots,d\text{.}$ Each component $g_{ij}$ of the local Riemannian metric $\bm{g}=(g_{ij})_{i,j=1}^{d}$ becomes
$g_{ij}=\delta_{ij}+\sum_{s=1}^{n-d}\frac{\partial q^{s}}{\partial\theta_{i}}\frac{\partial q^{s}}{\partial\theta_{j}}.$ (23)
Let the vector field $\bm{u}=\tilde{u}^{i}\partial_{i}$ in the local Monge coordinate system. Then the Bochner Laplacian takes the form
$\Delta_{B}\bm{u}(\mathbf{x}_{0})=g^{kj}\tilde{u}_{;kj}^{i}\partial_{i}|_{\mathbf{x}_{0}}=\sum_{k}\tilde{u}_{;kk}^{i}\partial_{i}|_{\mathbf{x}_{0}}=\sum_{k}\left(\partial_{kk}\tilde{u}^{i}+\tilde{u}^{l}\frac{\partial\Gamma_{kl}^{i}}{\partial\theta_{k}}\right)\partial_{i}|_{\mathbf{x}_{0}}.$
In the second equality, we use the fact that the metric $\bm{g}$ is the identity matrix at $\mathbf{x}_{0}$. By using the metric (23), the derivative of the Christoffel symbol at the base point can be calculated as,
$\quad\frac{\partial\Gamma_{kl}^{i}}{\partial\theta_{r}}\Big{|}_{\mathbf{x}_{0}}=\frac{1}{2}g^{im}\left(\frac{\partial g_{mk}}{\partial\theta_{r}\partial\theta_{l}}+\frac{\partial g_{ml}}{\partial\theta_{r}\partial\theta_{k}}-\frac{\partial g_{kl}}{\partial\theta_{r}\partial\theta_{m}}\right)\Big{|}_{\mathbf{x}_{0}}=\sum_{s=1}^{n-d}\partial_{ir}q^{s}\partial_{kl}q^{s}|_{\mathbf{x}_{0}},$ (24)
where $\partial_{kl}q^{s}=\frac{\partial^{2}q^{s}}{\partial\theta_{k}\partial\theta_{l}}$. Then the Bochner Laplacian can be written as
$\Delta_{B}\bm{u}(\mathbf{x}_{0})=\sum_{k}\left(\partial_{kk}\tilde{u}^{i}+\tilde{u}^{l}\sum_{s=1}^{n-d}\partial_{ik}q^{s}\partial_{kl}q^{s}\right)\partial_{i}|_{\mathbf{x}_{0}}.$ (25)
The results of the Hodge Laplacian and another vector Laplacian can be found in Section A.
3.3.2 Numerical implementation
Note that the formulation (25) requires computing the $n-d$ normal vectors $\left\{\bm{n}_{i}({\mathbf{x}_{0}})\right\}_{i=1}^{n-d}$ at the base point which are obtained using the QR decomposition algorithm with a computational cost of $O(n^{3})$. To avoid explicitly computing these normal vectors, we instead apply the GMLS to each component of ${\mathbf{x}}=(x^{1},\ldots,x^{n})$ in the stencil $S_{{\mathbf{x}_{0}}}=\left\{{\mathbf{x}}_{0,k}\right\}_{k=1}^{K}=\{x_{0,k}^{s}\}_{k=1,\ldots,K}^{s=1,\ldots,n}$ rather than using $q^{s}$ as defined in (22). For $s=1,\ldots,n$, the $s$th component of the Taylor expansion of $\mathbf{x}(\bm{\theta})$ with respect to $\bm{\theta}=(\theta_{1},...,\theta_{d})$ at $\bm{\theta}=\bm{0}$ is given by
$x^{s}=x_{0}^{s}+\sum_{i=1}^{d}t_{i}^{s}\theta_{i}+\sum_{2\leq\left|\bm{\alpha}\right|\leq l}b_{\alpha_{1},\ldots,\alpha_{d}}^{s}\theta_{1}^{\alpha_{1}}\cdots\theta_{d}^{\alpha_{d}}+O(\left\|\bm{\theta}\right\|^{l+1}),$ (26)
Here $t_{i}^{s}$ is the $s$th component of the tangent vectors $\bm{t}_{i}({\mathbf{x}_{0}})$ and $\mathbf{x}(\mathbf{0})=(x_{0}^{1},...,x_{0}^{n})$. We then define $p^{s}$ as
$p^{s}=x^{s}-x_{0}^{s}-\sum_{i=1}^{d}t_{i}^{s}\theta_{i}=\sum_{2\leq\left|\bm{\alpha}\right|\leq l}b_{\bm{\alpha}}^{s}\bm{\theta}^{\bm{\alpha}}.$ (27)
where $b_{\bm{\alpha}}^{s}$ will be determined by the GMLS approximation. For the terms where $|\bm{\alpha}|=2$ in (27), we can write them in the following quadratic form,
$\sum_{\left|\bm{\alpha}\right|=2}b_{\bm{\alpha}}^{s}\bm{\theta}^{\bm{\alpha}}=\left[\begin{array}[]{c}\theta_{1}\\ \vdots\\ \theta_{d}\end{array}\right]^{\top}\left[\begin{array}[]{ccc}c_{11}^{s}&\cdots&c_{1d}^{s}\\ \vdots&\ddots&\vdots\\ c_{d1}^{s}&\cdots&c_{dd}^{s}\end{array}\right]\left[\begin{array}[]{c}\theta_{1}\\ \vdots\\ \theta_{d}\end{array}\right],\text{ \ for }s=1,\ldots,n,$ (28)
where the coefficient matrix is symmetric with $c_{ii}^{s}=\{b_{\bm{\alpha}}^{s}|\alpha_{i}=2$ and the others $\alpha_{j}=0$ for $j\neq i\}$ and $c_{ik}^{s}=$$\{\frac{1}{2}b_{\bm{\alpha}}^{s}|\alpha_{i}=\alpha_{k}=1$ for $i\neq k$ and the others $\alpha_{j}=0$ for $j\neq i$ and $j\neq k\}$.
On the other hand, let the $s$th normal component $q^{s}$ in the Monge coordinate (22) be
$q^{s}=\sum_{2\leq\left|\bm{\alpha}\right|\leq l}a_{\bm{\alpha}}^{s}\bm{\theta}^{\bm{\alpha}},\text{ \ for }s=1,\ldots,n-d,$ (29)
which generalizes equation (10) from dimension $2$ to dimension $d$. Then we have the following proposition with its proof provided in Section A.3.
Proposition 3.2.
Let $\left\{p^{i}\right\}_{i=1}^{n}$ and $\left\{q^{j}\right\}_{j=1}^{n-d}$ be defined as in (27) and (29), respectively. Then for any integers $1\leq i,r,k,l\leq d,$
$\sum_{s=1}^{n-d}\partial_{ir}q^{s}\partial_{kl}q^{s}=\sum_{s=1}^{n}\partial_{ir}p^{s}\partial_{kl}p^{s},$ (30)
where the derivative $\partial_{ir}=\frac{\partial^{2}}{\partial\theta_{i}\partial\theta_{r}}$ is with respect to $\left\{\theta_{i}\right\}_{i=1}^{d}$. Moreover, at the base point $\mathbf{x}_{0}$, it holds that
$\sum_{s=1}^{n-d}\partial_{ir}q^{s}\partial_{kl}q^{s}|_{\mathbf{x}_{0}}=\sum_{s=1}^{n}4c_{ir}^{s}c_{kl}^{s},$ (31)
where $c_{ir}^{s}$ is defined in (28).
By using the equality (31), the Bochner Laplacian in (25) can be written as
$\Delta_{B}\bm{u}(\mathbf{x}_{0})=\sum_{k}\left(\partial_{kk}\tilde{u}^{i}+\tilde{u}^{l}\sum_{s=1}^{n}4c_{ir}^{s}c_{kl}^{s}\right)\partial_{i}|_{\mathbf{x}_{0}}.$ (32)
We note that the Bochner Laplacian involves only the coefficients of the GMLS approximation of $p^{s}$, corresponding to ${|\bm{\alpha}|=2}$ in (28).
3.3.3 Algorithm and complexity analysis
@@
The complete procedure of the intrinsic GMLS approximation of the Bochner Laplacian (in the case when $d\ll n$) and its complexity analysis are outlined in Algorithm 1. Notably, one can follow the similar procedure to approximate other vector Laplacians acting on vector fields using the intrinsic method (See Section A).
Algorithm 1 Intrinsic GMLS of the Bochner Laplacian when $d\ll n$
@@
1:  Input: A point cloud $\{\mathbf{x}_{i}\}_{i=1}^{N}\subset M$, bases of (estimated) tangent vectors at each node $\left\{\bm{t}_{1}(\mathbf{x}_{i}),\ldots,\bm{t}_{d}(\mathbf{x}_{i})\right\}_{i=1}^{N}$, the degree $l$ of polynomials, and a parameter $K(>m)$ nearest neighbors where $m=\left(\begin{array}[]{c}l+d\\ d\end{array}\right)$ is the number of monomial basis functions $\left\{\bm{\theta}^{\bm{\alpha}}|0\leq\left|\bm{\alpha}\right|\leq l\right\}$.
2:  Set $\mathbf{L}_{B}$ to be a sparse $dN\times dN$ matrix with $d^{2}NK$ nonzeros.
@@
5:     Find the GMLS approximation of $p^{s}=x^{s}-x_{0}^{s}-\sum_{i=1}^{d}t_{i}^{s}\theta_{i}$ as in (27) using the optimal solution $\hat{p}^{s}=\sum_{2\leq\left|\bm{\alpha}\right|\leq l}b_{\bm{\alpha}}^{s}\bm{\theta}^{\bm{\alpha}}$ to the least-squares problem with the complexity of $O(NnKm^{2}+Nnm^{3})$.
6:     Construct the local tangent-vector bases $\{\partial_{j}|_{\mathbf{x}_{i,k}}\}_{j=1,\ldots,d}$ over $S_{\mathbf{x}_{i}}$ with the complexity $O(NnKm^{2}d)$ based on taking the derivative of (26).
7:     Compute the coefficients of the Bochner Laplacian $\{\bm{\tilde{w}}^{k}\}_{k=1}^{K}$ in the local Monge coordinate based on (32) with the complexity $O(Nnd^{3})$.
8:     Calculate the weights $\{\bm{w}^{k}\}_{k=1}^{K}$ with respect to the global tangent-vector bases with the complexity $O(NKd^{2}n+NKd^{4})$ by the formula,
$\bm{w}^{k}=\bm{\tilde{w}}^{k}\left(\left[\begin{array}[]{c}\partial_{1}^{\top}|_{{\mathbf{x}_{0,k}}}\\ \vdots\\ \partial_{d}^{\top}|_{{\mathbf{x}_{0,k}}}\end{array}\right]\left[\begin{array}[]{ccc}\partial_{1}|_{{\mathbf{x}_{0,k}}}&\cdots&\partial_{d}|_{{\mathbf{x}_{0,k}}}\end{array}\right]\right)^{-1}\left(\left[\begin{array}[]{c}\partial_{1}^{\top}|_{{\mathbf{x}_{0,k}}}\\ \vdots\\ \partial_{d}^{\top}|_{{\mathbf{x}_{0,k}}}\end{array}\right]\left[\begin{array}[]{ccc}\bm{t}_{1}({\mathbf{x}_{0,k}})&\cdots&\bm{t}_{d}({\mathbf{\ x}_{0,k}})\end{array}\right]\right).$
9:     Arrange the weights $\{\bm{w}^{k}\}_{k=1}^{K}$ into corresponding rows and columns of $\mathbf{L}_{B}$.
10:  end for
11:  Output: The approximate operator matrix $\mathbf{L}_{B}$.
@@ 3.4 Approximation of covariant derivative
Note that the covariant derivative of the vector field $\bm{u}=\tilde{u}^{i}\partial_{i}$ along itself could be written as
$\nabla_{\bm{u}}\bm{u}=\tilde{u}_{;k}^{i}\tilde{u}^{k}\partial_{i}=\tilde{u}^{k}\left(\partial_{k}\tilde{u}^{i}+\tilde{u}^{j}\Gamma_{kj}^{i}\right)\partial_{i}.$
By using the local Monge coordinate system introduced in Section 3.1, the approximation of $\nabla_{\bm{u}}\bm{u}$ on 2D manifolds could be derived directly as
$\nabla_{\bm{u}}\bm{u}(\mathbf{x}_{0})$ $=$ $\tilde{u}^{k}\left(\partial_{k}\tilde{u}^{i}+\tilde{u}^{j}\Gamma_{kj}^{i}\right)\partial_{i}\big{|}_{\mathbf{x}_{0}}$
@@
where the coefficients $\beta_{\alpha_{1},\alpha_{2}}$ and $\gamma_{\alpha_{1},\alpha_{2}}$ for $0\leq\alpha_{1}+\alpha_{2}\leq l$ are defined in (18). Here, we again use Proposition 3.1 to simplify the calculations. We remark that this formulation naturally extends to higher dimensions.